OpenStudy (anonymous):

-Medal will be given- A gardener found that he was able to plant ¼ of a packet of flower seeds in 1/5 of a garden. At this rate, how much of the garden would he cover with the entire packet of seeds? A. 1/20 B. 4/5 C. 5/4 D. 9/2

OpenStudy (anonymous):

I think it is B., but my teacher says it is C. I think he is tricking us, but I can't be sure.

OpenStudy (texaschic101):

(1/4)/(1/5) = (4/4)/x cross mutiply... 1/4x = (1/5)(4/4) 1/4x = 4/20 x = (4/20)/(1/4) x = 4/20 x 4/1 x = 16/20 = 4/5 I am going to agree with you @justin12lo ....I think it is 4/5 (B) :)

OpenStudy (texaschic101):

@jim_thompson5910 ....can you check this....please

OpenStudy (texaschic101):

@hartnn ...can you check this pease

OpenStudy (texaschic101):

is the teacher wrong ?

OpenStudy (texaschic101):

@Mertsj ....can you check this

jimthompson5910 (jim_thompson5910):

1/4 packet = 1/5 garden 4*(1/4) packet = 4*(1/5) garden 4/4 packet = 4/5 garden 1 packet = 4/5 garden

jimthompson5910 (jim_thompson5910):

so I agree, it's B

OpenStudy (texaschic101):

I just set it up as a proportion

OpenStudy (texaschic101):

thank you so much

OpenStudy (mertsj):